OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. (4 points) Part 2: Provide the solution to the system. (2 points) 5x – 9y = –16 2x + 6y = –16

OpenStudy (anonymous):

i would multiply the first equations by \(2\) and the second equation by \(3\) that would give \[10x-18y=-32\]\[6x+18y=-48\] then i would add the two equations together to get \[16x=-80\] divide by \(16\) to get \(x=-5\)

OpenStudy (anonymous):

Thank u!!

OpenStudy (anonymous):

now that you know \(x=-5\) you can substitute \(-5\) for \(x\) in either equation and solve for \(y\)

OpenStudy (anonymous):

\[2(-5)+6y=-16\]\[-10+6y=-16\]\[6y=-6\]\[y=-1\]

OpenStudy (anonymous):

yw